OpenStudy (anonymous):

What is the sum of a 22-term arithmetic sequence where the first term is 54 and the last term is -30?

OpenStudy (anonymous):

as before here reqd S= 22/2 * (54-30) =11*24 =264

OpenStudy (sirm3d):

use the formula \[\Large S_n=\frac{n}{2}\left(t_1 + t_n\right)\] where\(S_n\) is the sum of the \(n\) terms whose \(1^\text {st}\) and \(n^\text {th}\) terms are \(t_1\) and \(t_n\)